A sample of 22 adult elephants has a mean weight of 14,500 lbs with a sample standard deviation of 1000lbs. Find the talpha/2 for a 95% confidence interval of the true mean.
Answer by solve_for_x(190)

To find t(alpha/2), you need to know the number of degrees of freedom, and the value of alpha.

For a 95% confidence interval, the value of alpha is 100% - 95% = 5% = 0.05.

alpha/2 is then 0.05/2 = 0.025

The number of degrees of freedom for a t-test is one less than the sample size:

df = n - 1 = 22 - 1 = 21

Then, using a table of the t-distribution for alpha = 0.05 or alpha/2 = 0.025 (depending
on how the table is set up), move down to the row for 21 degrees of freedom and read off
the value:

t(alpha/2) = 2.080

You can also get this value from Excel by entering "=T.INV.2T(0.05,21))" into a cell.
Note that this function returns a two-tailed value, so you must use the full value of
alpha instead of alpha/2.